If a coin is tossed thrice, find the probability of a getting head at least two times.


1
Expert's answer
2022-05-23T14:39:46-0400

Let's look into the possible outcomes.

Sample space: {HHH, HTH,THH,TTH, HHT, HTT,THT,TTT​}

Total number of possible outcomes=8

Probability of getting at least two heads is:

P(A)=P(getting two heads)+P(getting 3 heads)

= 3/8 + 1/8

= 4/8 = 1/2